# Split Two-Periodic Aztec Diamond

### v3

@@ 3. Derivation of the Correlation Kernel
Our main goal in this section is to prove the results stated in Theorem 2.1. In order to prove Theorem 2.1, we start in Section 3.1 by stating an intermediate version of the kernel which can be derived by using techniques from [3]. The form of the kernel differs slightly depending on whether the coordinate $(4m^{\prime},2\xi^{\prime}+j)$ is on the $\alpha$-side or $\beta$-side of the interface. In Section 3.2, we prove the statement of the kernel for $m^{\prime}\leq\nicefrac{{N}}{{2}}$ and in Section 3.3, we prove the statement of the kernel for $m^{\prime}>\nicefrac{{N}}{{2}}$. After deriving the kernel, we prove the statement given in Corollary 2.1, which illustrates the symmetry of the kernel and the relationship of the kernel to the typical two-periodic Aztec diamond. This work is done in Section 3.4.
@@ 3.1. An Intermediate Correlation Kernel
Applying this double contour integral into the formula in (56), yields the expression presented in equation (16).
3.4. Proof of Corollary 2.1
We have two goals in mind for wanting to prove Corollary 2.1. Firstly, the kernels on the $\alpha$-side and $\beta$-side of the interface now have a symmetry to them and we are free to limit our computations by looking at a single side of the model. [^7] In the case of the typical two-periodic Aztec diamond, this kernel symmetry is obvious by flipping the Aztec diamond twice (vertically and horizontally) and then gauge transforming the edge weights. In the case of the split two-periodic model, you can no longer use a gauge transformation to get back to the original edge weights as the interface causes a problem. Secondly, the adjusted form of the kernel contains the correlation kernel for the typically two-periodic Aztec diamond, so we will more easily be able to make comparisons.
The main process we take in this proof is move the $w$-contour of the double contour integral to the branch cuts and then through the branch cuts on to the second sheet of the Riemann surface that was described earlier in Section 2.3.
Proof.
We start with the form of the kernel presented in equation (16). We make the following observations about the $F$-matrices,
(60) $\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)=F_{\beta,1}(w)-\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,2}(w)F_{\beta,2}(w)$
Next, we substitute this equality into the double contour integral in equation (16). We write the result as two separate contour integrals,
(61) $\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ =\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}F_{\beta,1}(w)\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ -\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(w)F_{\alpha,2}(w)F_{\beta,2}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
The first double contour integral matches up with the first double contour integral in (15), so we do not need to manipulate it further.
We set the branch cuts of the $w$ part of the second double contour integral to be along the negative real axis. Let $\delta=\max\{\alpha,\beta\}$, then the branch cuts are $(-\infty,-\delta^{-2}]\cup[-\delta^{2},0]$. We deform the contour $\gamma_{1}$ out to infinity and around, so that the new contour surrounds only the branch cuts. From this maneuver, we get an additional single contour integral residue at $z=w$. This single contour integral has the form,
$\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,1}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(z)F_{\alpha,2}(z)F_{\beta,2}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
Next we will pass the $w$ contour of the double contour integral through the branch cuts and on to the second sheet of the Riemann surface. Pushing the contour to the other sheet means we need to make the following changes to the integral: $r_{\beta,1}(w)\to r_{\beta,2}(w)$ and
$\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,2}(w)F_{\beta,2}(w)\to\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,2}(w)F_{\alpha,1}(w)F_{\beta,1}(w)$
These are the analytic continuations of the functions on the other sheet. Also, pushing the contour through to the second sheet changes the direction of the contour, introducing a minus sign. We have thus shown,
(62) $-\oint_{\gamma_{1}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(w)F_{\alpha,2}(w)F_{\beta,2}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ =\oint_{\gamma_{\text{Br}}}\oint_{\gamma_{0,1}}\frac{\text{d}w\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,2}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,2}(w)F_{\alpha,1}(w)F_{\beta,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{2\pi\textrm{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi^{\prime}-\xi}r_{\beta,1}(z)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(z)F_{\alpha,2}(z)F_{\beta,2}(z)}{1+2g_{\alpha,\beta}(z)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
Where $\gamma_{\text{Br}}$ denotes the contour surround the branch cuts (in the clockwise direction). ∎
@@ 4. Local Asymptotics of the Model
The goal of this section is to prove the local asymptotic results stated in Propositions 2.2 and 2.3. We begin in Section 4.1 by stating some necessary results about the behavior of the saddle functions. In particular, we need precise statements about the location of saddle points and the contours of steepest ascent and descent in the smooth and rough regions of the model. Once we have the necessary preliminary information, we prove Propositions 2.1-2.3 in Sections 4.2, 4.3, and 4.4.
@@ 4.1. Preliminaries of Saddle Functions
where $\psi_{\varepsilon,k}(z)$ is defined in equation (22). We start by stating a basic lemma regarding the asymptotic behavior of the function $\operatorname{Re}\psi_{k}(z)$.
@@
    if $(x,y)\in\mathfrak{S}$, then $z_{1}^{*}$ is a local maximum, otherwise it is a local minimum.
@@
(64) $\lim_{z\to(-\alpha^{2})^{-}}\frac{r^{\prime}_{\alpha,1}(z)}{r_{\alpha,1}(z)}=+\infty\qquad\lim_{z\to(-\alpha^{-2})^{+}}\frac{r^{\prime}_{\alpha,1}(z)}{r_{\alpha,1}(z)}=-\infty$
(65) $\lim_{z\to(-\alpha^{2})^{-}}\frac{r^{\prime}_{\alpha,2}(z)}{r_{\alpha,2}(z)}=-\infty\qquad\lim_{z\to(-\alpha^{-2})^{+}}\frac{r^{\prime}_{\alpha,2}(z)}{r_{\alpha,2}(z)}=+\infty$
From this we deduce that $z_{2}^{*}$ is a local maximum of $\operatorname{Re}\psi_{2}(z)$ and $z_{1}^{*}$ is a local minimum of $\operatorname{Re}\psi_{1}(z)$. ∎
The contour of steepest ascent from $z_{k}^{*}$ is the contour through $z_{k}^{*}$ where $\operatorname{Im}\psi_{k}(z)$ is held constant and $\operatorname{Re}\psi_{k}(z)$ increases the fastest as we move away from the saddle point. Since $z_{2}^{*}$ is a maximum of $\operatorname{Re}\psi_{2}(z)$, the path of steepest ascent is perpendicular to the real axis at $z_{2}^{*}$ (the same is true for $z_{1}^{*}$ when the asymptotic coordinate is in the smooth region). We define the following regions of the plane, [^9] While this notation is the same as in [11] the regions will be reversed in our notation due to the change in sign of the saddle functions presented here.
@@
    $\Omega^{+}_{2}$ is a bounded set containing a single component which contains $\gamma_{s,2}\backslash\{z_{2}^{*}\}$.
@@
  1. (i)
  2. (ii)
    $\Omega^{+}_{1}$ is a bounded set containing at most three components which contains $\gamma_{s,1}\backslash\{z_{1}^{*}\}$.
  3. (iii)
Figure 9 shows an example of the regions $\Omega_{2}^{+}$ and $\Omega_{2}^{-}$ along with the steepest ascent contour. This lemma is a slightly modified version of Lemma 6.9 of [11]. The reader should refer to their text for a the proof. Note that the proof given also holds for $\psi_{2}(z)$ as long as the asymptotic coordinate of the model is to the left of the interface (and regardless of the macroscopic region of the model). In the first part, we conclude that $\Omega^{+}_{2}$ is a single component, because there is exactly one saddle point on the interval $(-\alpha^{-2},-\alpha^{2})$.
Figure 9. An example of the steepest ascent contour of $\psi_{2}(z;x,y)$ through $z_{2}^{*}$ and the regions $\Omega^{+}_{2}$ and $\Omega^{-}_{2}$. The steepest ascent contour is shown in red, $\Omega^{+}_{2}$ is the region in yellow, and $\Omega^{-}_{2}$ is the region in blue. Contour plot was generated with Mathematica.
The above information is enough for the analysis of $I_{2,2}^{\alpha}$ and for the analysis of $I^{\alpha}_{2,1}$ in the smooth region. In order to analyze $I^{\alpha}_{2,1}$ in the rough region, we will need some information on the contours of steepest descent and ascent in the rough region. For $x\in(0,\nicefrac{{1}}{{2}})$ and $(x,y)\in\mathfrak{R}$, the saddle function $\psi_{1}(z)$ has three saddle point, $z_{1}^{*}$, $z_{3}^{*}$, and $z_{4}^{*}$, where the latter two are complex. In particular, $z_{3}^{*}=\overline{z_{4}^{*}}$. Let $z_{3}^{*}$ denote the complex saddle point with positive imaginary part. We define the following regions of the complex plane,
(68) $\Pi^{+}=\{z\in\mathbb{C}:\operatorname{Re}\psi_{1}(z)>\operatorname{Re}\psi_{1}(z_{3}^{*})\}$
(69) $\Pi^{-}=\{z\in\mathbb{C}:\operatorname{Re}\psi_{1}(z)<\operatorname{Re}\psi_{1}(z_{3}^{*})\}$
It suffices to define these regions just using $z_{3}^{*}$, since $\operatorname{Re}\psi_{1}(z_{3}^{*})=\operatorname{Re}\psi_{1}(z_{4}^{*})$. We now describe some important aspects about these regions,
@@
We will focus only on the upper half plane, since the regions $\Pi^{\pm}$ are symmetric about the real axis. The boundary between these regions are contours where $\operatorname{Re}\psi_{1}(z)=\operatorname{Re}\psi_{1}(z_{3}^{*})$. These contours remain bounded, since $\operatorname{Re}\psi_{1}(z)\to-\infty$ as $z\to\infty$. Additionally, these contours do not intersect anywhere in the upper half plane or on the real axis, because this would violate the maximum principle of harmonic functions. This means that the four half-paths emanating from $z_{3}^{*}$ intersect the real axis at four different points. To see the possible intersection points, we note the following limits,
  • $\lim_{z\to\infty}\operatorname{Re}\psi_{1}(z)=-\infty$
  • $\lim_{z\to 0}\operatorname{Re}\psi_{1}(z)=-\infty$
  • $\lim_{z\to 1}\operatorname{Re}\psi_{1}(z)=+\infty$
@@
  • $\operatorname{Re}\psi_{1}(z)$ is strictly decreasing on $(1,\infty)$
  • $\operatorname{Re}\psi_{1}(z)$ is strictly increasing on $(0,1)$
  • $\operatorname{Re}\psi_{1}(z)$ is strictly increasing on $(-\infty,-\alpha^{-2})$
  • $\operatorname{Re}\psi_{1}(z)$ is strictly decreasing on $(-\alpha^{2},z_{1}^{*})$
  • $\operatorname{Re}\psi_{1}(z)$ is strictly increasing on $(z_{1}^{*},-\alpha^{2})$
  • $\operatorname{Re}\psi_{1}(z)$ is strictly decreasing on $(-\alpha^{2},0)$
@@
Since $\operatorname{Re}\psi_{1}(z)$ is not strictly increasing on the interval $(-\infty,0)$, where the two half-paths intersect this interval can vary. Either $-\alpha^{2}$ or $-\alpha^{-2}$ is a global maximum of $\operatorname{Re}\psi_{1}(z)$ on $(-\infty,0)$, so at least one of them (possibly both) must be contained in $\Pi^{+}$. Alternatively, one of the branch points may be contained in a separate third component that intersect the branch cut. [^10] It is unclear whether this actually occurs, but it does not affect our analysis either way. There can only be one additional component since any additional components must intersect the branch cuts and $\operatorname{Re}\psi_{1}(x)$ is monotonic along the branch cuts. Similarly, there is no additional component of $\Pi^{-}$ because it would have to be away from the branch cuts and thus it would break the maximum/minimum principle of harmonic functions. ∎
@@
Figure 11. Steepest descent and ascent contours for $\psi_{1}(z)$ in the rough region. The point where the contours cross are the two complex saddle points. The red contour is the path of steepest descent, while the blue contour is the path of steepest ascent.
4.2. Proof of Proposition 2.1
As stated below equations (24) and (25), we will assume the coordinates $(m,\xi)$ and $(m^{\prime},\xi^{\prime})$ are on the same side of the interface. Regardless of which side they are on, $J$ has the form,

### v4

@@ 3. Derivation of the Correlation Kernel
Our main goal in this section is to prove the results stated in Theorem 2.1. In order to prove Theorem 2.1, we start in Section 3.1 by stating an intermediate version of the kernel which can be derived by using techniques from [3]. The form of the kernel differs slightly depending on whether the coordinate $(4m^{\prime},2\xi^{\prime}+j)$ is on the $\alpha$-side or $\beta$-side of the interface. In Section 3.2, we prove the statement of the kernel for $m^{\prime}\leq\nicefrac{{N}}{{2}}$ and in Section 3.3, we prove the statement of the kernel for $m^{\prime}>\nicefrac{{N}}{{2}}$.
@@ 3.1. An Intermediate Correlation Kernel
We now have the following form of the correlation kernel,
(57) $\Big{[}\mathbb{K}_{N}(4m^{\prime},2\xi^{\prime}+j;4m,2\xi+i)\Big{]}_{i,j=0}^{1}=-\frac{\mathbb{I}_{m<m^{\prime}}}{2\pi\emph{i}}\oint_{\gamma_{0,1}}\frac{\text{d}z}{z}z^{\xi-\xi^{\prime}}\Phi_{\beta}(z)^{\frac{N}{2}-m^{\prime}}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}\\ +\frac{1}{(2\pi\emph{i})^{2}}\oint_{\gamma_{1}}\text{d}w\oint_{\gamma_{0,1}}\frac{\text{d}z}{z(z-w)}\frac{w^{\xi^{\prime}+N}(z-1)^{N}}{z^{\xi+N}(w-1)^{N}}r_{\beta,1}(w)^{\frac{N}{2}-m^{\prime}}\frac{2F_{\beta,1}(w)F_{\alpha,1}(w)}{1+2g_{\alpha,\beta}(w)}\Phi_{\varepsilon}(z)^{m-\frac{N}{2}}$
for $m^{\prime}>\nicefrac{{N}}{{2}}$. This is a valid kernel that is amenable to asymptotic analysis, however we wish to make one additional change to the kernel so that it can better be compared to the case of $m^{\prime}\leq\nicefrac{{N}}{{2}}$. To do so, we use the following identity,
$\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\beta,1}(w)F_{\alpha,1}(w)=F_{\alpha,1}(w)-\frac{2}{1+2g_{\alpha,\beta}(w)}F_{\alpha,1}(w)F_{\beta,2}(w)F_{\alpha,2}(w)$
The result of this substitution is the kernel presented in (16).
@@ 4. Local Asymptotics of the Model
The goal of this section is to prove the local asymptotic results stated in Propositions 2.1 and 2.2. We begin in Section 4.1 by stating some necessary results about the behavior of the saddle functions. In particular, we need precise statements about the location of saddle points and the contours of steepest ascent and descent in the smooth and rough regions of the model. Once we have the necessary preliminary information, we prove Propositions 2.1 and 2.2 in Sections 4.3 and 4.4, respectively.
@@ 4.1. Preliminaries of Saddle Functions
where $\psi_{\varepsilon,k}(z)$ is restated below.
$\psi_{\varepsilon,k}(z)=(1+y)\log z-\log(z-1)+\left(\frac{1}{2}-x\right)\log r_{\varepsilon,k}(z)$
We start by stating a basic lemma regarding the asymptotic behavior of the function $\operatorname{Re}\psi_{k}(z)$.
@@
    if $(x,y)\in\mathfrak{S}$, then $z_{2}^{*}$ is a local maximum of $\psi_{2}(z)$, otherwise it is a local minimum.
@@
(59) $\lim_{z\to(-\alpha^{2})^{-}}\frac{r^{\prime}_{\alpha,1}(z)}{r_{\alpha,1}(z)}=-\infty\qquad\lim_{z\to(-\alpha^{-2})^{+}}\frac{r^{\prime}_{\alpha,1}(z)}{r_{\alpha,1}(z)}=+\infty$
(60) $\lim_{z\to(-\alpha^{2})^{-}}\frac{r^{\prime}_{\alpha,2}(z)}{r_{\alpha,2}(z)}=+\infty\qquad\lim_{z\to(-\alpha^{-2})^{+}}\frac{r^{\prime}_{\alpha,2}(z)}{r_{\alpha,2}(z)}=-\infty$
From this we deduce that $z_{2}^{*}$ is a local minimum of $\operatorname{Re}\psi_{2}(z)$ and $z_{1}^{*}$ is a local maximum of $\operatorname{Re}\psi_{1}(z)$. ∎
Since we are restricting our analysis to $x<\nicefrac{{1}}{{2}}$ we can make some additional remarks regarding the general location of the saddle points. Under this restriction, $\psi_{1}(z)$ always has a single saddle point, while $\psi_{2}(z)$ has three saddle points (unless the point $(x,y)$ is on a macroscopic boundary). We will always refer to the other two saddle points of $\psi_{2}(z)$ as $z_{3}^{*}$ and $z_{4}^{*}$. When $(x,y)\in\mathfrak{S}$, these additional saddle points also lie on the interval $(-\alpha^{-2},-\alpha^{2})$. We adopt the convention, $z_{3}^{*}<z_{2}^{*}<z_{4}^{*}$. The points $z_{3}^{*}$ and $z_{4}^{*}$ are both local minimums of $\operatorname{Re}\psi_{2}(z)$.
The contour of steepest ascent from $z_{k}^{*}$ is the contour through $z_{k}^{*}$ where $\operatorname{Im}\psi_{k}(z)$ is held constant and $\operatorname{Re}\psi_{k}(z)$ increases the fastest as we move away from the saddle point. Since $z_{1}^{*}$ is a maximum of $\operatorname{Re}\psi_{1}(z)$, the path of steepest ascent is perpendicular to the real axis at $z_{1}^{*}$ (the same is true for $z_{2}^{*}$ when the asymptotic coordinate is in the smooth region). We define the following regions of the complex plane, [^9] While this notation is the same as in [11] the regions will be reversed in our notation due to the change in sign of the saddle functions presented here.
@@
  2. (ii)
    $\Omega^{+}_{1}$ is a bounded set containing a single component which contains $\gamma_{s,1}\backslash\{z_{1}^{*}\}$.
  3. (iii)
  1. (i)
@@
    $\Omega^{+}_{2}$ is a bounded set containing at most three components that are a positive distance apart. One of which contains $\gamma_{s,2}\backslash\{z_{2}^{*}\}$.
@@
Figure 9 shows an example of the regions $\Omega_{1}^{+}$ and $\Omega_{1}^{-}$ along with the steepest ascent contour. This lemma is a slightly modified version of Lemma 6.9 of [11]. The reader should refer to their text for a the proof. One can check that the proof is valid for $\psi_{1}(z)$ for any $x<\nicefrac{{1}}{{2}}$. This is because $z_{1}^{*}$ is always a local maximum of $\operatorname{Re}\psi_{1}(z)$. In the first part, we conclude that $\Omega^{+}_{1}$ is a single component, because there is exactly one saddle point on the interval $(-\alpha^{-2},-\alpha^{2})$. The case of three components is only possible if there are three saddle points on the interval $(-\alpha^{-2},-\alpha^{2})$.
In our asymptotic proofs we will use a statement a bit stronger than (both) statement (iii) above. In particular, we state the following corollary,
Corollary 4.1.
[…]
Proof.
In the case where $\Omega_{j}^{+}$ is only one component, the statement immediately follows from the fact that $z_{1}^{*}$ is the only point on the negative not a positive distance away from $\Omega^{+}_{j}$. In the case where $\Omega_{j}^{+}$ has three components, we must also invoke the fact that the components are all a positive distance away from each other. ∎
Figure 9. An example of the steepest ascent contour of $\psi_{1}(z;x,y)$ through $z_{1}^{*}$ and the regions $\Omega^{+}_{1}$ and $\Omega^{-}_{1}$. The steepest ascent contour is shown in red, $\Omega^{+}_{1}$ is the region in yellow, and $\Omega^{-}_{1}$ is the region in blue. Contour plot was generated with Mathematica.
The above information is enough for the analysis of $I_{2,1}^{\alpha}$ and for the analysis of $I^{\alpha}_{2,2}$ in the smooth region. In order to analyze $I^{\alpha}_{2,2}$ in the rough region, we will need some information on the contours of steepest descent and ascent in the rough region. For $x\in(0,\nicefrac{{1}}{{2}})$ and $(x,y)\in\mathfrak{R}$, the additional saddle points of $\psi_{2}9z)$, $z_{3}^{*}$ and $z_{4}^{*}$, are complex. In particular, $z_{3}^{*}=\overline{z_{4}^{*}}$. We will let $z_{3}^{*}$ denote the complex saddle point with positive imaginary part. We define the following regions of the complex plane,
(63) $\Pi^{+}=\{z\in\mathbb{C}:\operatorname{Re}\psi_{2}(z)>\operatorname{Re}\psi_{2}(z_{3}^{*})\}$
(64) $\Pi^{-}=\{z\in\mathbb{C}:\operatorname{Re}\psi_{2}(z)<\operatorname{Re}\psi_{2}(z_{3}^{*})\}$
It suffices to define these regions just using $z_{3}^{*}$, since $\operatorname{Re}\psi_{2}(z_{3}^{*})=\operatorname{Re}\psi_{2}(z_{4}^{*})$. We now describe some important aspects about these regions,
@@
We will focus only on the upper half plane, since the regions $\Pi^{\pm}$ are symmetric about the real axis. The boundary between these regions are contours where $\operatorname{Re}\psi_{2}(z)=\operatorname{Re}\psi_{2}(z_{3}^{*})$. These contours remain bounded, since $\operatorname{Re}\psi_{2}(z)\to-\infty$ as $z\to\infty$. Additionally, these contours do not intersect anywhere in the upper half plane or on the real axis, because this would violate the maximum principle of harmonic functions. This means that the four half-paths emanating from $z_{3}^{*}$ intersect the real axis at four different points. To see the possible intersection points, we note the following limits,
  • $\lim_{z\to\infty}\operatorname{Re}\psi_{2}(z)=-\infty$
  • $\lim_{z\to 0}\operatorname{Re}\psi_{2}(z)=-\infty$
  • $\lim_{z\to 1}\operatorname{Re}\psi_{2}(z)=+\infty$
@@
  • $\operatorname{Re}\psi_{2}(z)$ is strictly decreasing on $(1,\infty)$
  • $\operatorname{Re}\psi_{2}(z)$ is strictly increasing on $(0,1)$
  • $\operatorname{Re}\psi_{2}(z)$ is strictly increasing on $(-\infty,-\alpha^{-2})$
  • $\operatorname{Re}\psi_{2}(z)$ is strictly decreasing on $(-\alpha^{2},z_{1}^{*})$
  • $\operatorname{Re}\psi_{2}(z)$ is strictly increasing on $(z_{1}^{*},-\alpha^{2})$
  • $\operatorname{Re}\psi_{2}(z)$ is strictly decreasing on $(-\alpha^{2},0)$
@@
Since $\operatorname{Re}\psi_{2}(z)$ is not strictly increasing on the interval $(-\infty,0)$, where the two half-paths intersect this interval can vary. Either $-\alpha^{2}$ or $-\alpha^{-2}$ is a global maximum of $\operatorname{Re}\psi_{2}(z)$ on $(-\infty,0)$, so at least one of them (possibly both) must be contained in $\Pi^{+}$. Alternatively, one of the branch points may be contained in a separate third component that intersect the branch cut. [^10] It is unclear whether this actually occurs, but it does not affect our analysis either way. There can only be one additional component since any additional components must intersect the branch cuts and $\operatorname{Re}\psi_{2}(x)$ is monotonic along the branch cuts. Similarly, there is no additional component of $\Pi^{-}$ because it would have to be away from the branch cuts and thus it would break the maximum/minimum principle of harmonic functions. ∎
@@
Figure 11. Steepest descent and ascent contours for $\psi_{2}(z)$ in the rough region. The point where the contours cross are the two complex saddle points. The red contour is the path of steepest descent, while the blue contour is the path of steepest ascent.
4.2. Initial Deformation of the Contours of $I_{2,k}^{\alpha}$
Before we detail the main asymptotic analysis of the contour integrals $I^{\alpha}_{2,k}$, we will first deform the $w$-contours of the integral from whats presented in equation (25). We first push the $w$-contour, $\gamma_{1}$, out to infinity. This maneuver is allowed since there is no singularity at $w=\infty$. Then we wrap the contour around the branch cuts. The exact branch cuts depend on the relative magnitudes of $\alpha$ and $\beta$ [^11] If $10<\alpha<\beta<$ 1 the branch cuts are $(-\infty,-\beta^{-2}]\cup[-\beta^{2},0]$, otherwise if $10<\beta<\alpha<$ 1 they are $(-\infty,-\alpha^{-2}]\cup[-\alpha^{2},0]$. We will denote this new contour $\gamma_{\text{Br}}$. The new contours are shown in Figure 12.
$\gamma_{\text{Br},1}$$\gamma_{\text{Br},2}$$\gamma_{0,1}$
Figure 12. Depiction of the new contours described in Section 4.2, where $\gamma_{\text{Br}}=\gamma_{\text{Br},1}\cup\gamma_{\text{Br},2}$. The contours $\gamma_{\text{Br},1}$ and $\gamma_{\text{Br},2}$ are oriented clockwise.
Since the $w$-contour passes over $\gamma_{0,1}$ when performing this procedure, it introduces an extra contour integral because of the residue at $w=z$. So we have,